# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient DRAM Maintenance Operations

### v8

@@ 1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4, 5, 6]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrinks, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [7, 8, 9, 10, 11, 12, 13, 14, 2, 15, 5, 16, 17, 18]. A modern DRAM chip requires three major types of maintenance operations (described in detail in section 5) for reliable and secure operation: 1) DRAM refresh [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], 2) RowHammer protection [47, 48, 49, 50, 12, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 18, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122], and 3) memory scrubbing [123, 124, 125, 126, 127, 128, 129, 130, 131, 17, 132, 133, 134, 135].111Memory scrubbing is not always required in consumer systems but often used in cloud systems. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [119, 136, 137]) and introducing new types of maintenance operations (e.g., targeted refresh [66, 64, 138], DDR5 RFM [119], and PRAC [119] as RowHammer defenses).222 A very recent update to the DDR5 standard [119] introduces PRAC, which is an on-DRAM-die read disturbance mitigation mechanism. PRAC requires more changes to the DRAM interface and continues to use RFM. Note that PRAC is concurrent with this work, as the initial version of this paper [139] was placed on arXiv on 27 July 2022 and initial submission to the MICRO 2022 conference was made on 22 April 2022.
@@
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.444SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors [4, 6] because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers). To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section 4.4). While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations as SMD overlaps the latency of a maintenance operation in a DRAM region with memory accesses to other DRAM regions.
@@
To enable practical adoption of SMD, we propose to implement it with low-cost modifications to existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section 6). Second, SMD implements a small (e.g., 16 bits per DRAM bank) memory structure called the Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [148, 149, 19]). In section 6, we show that these modifications have low DRAM chip area (1.1% of a DRAM chip) and negligible latency overhead (only 0.4% increase in row activation latency).
@@
We demonstrate the usefulness and versatility of SMD by implementing three in-DRAM maintenance mechanisms for DRAM refresh (section 5.1), RowHammer protection (section 5.2), and memory scrubbing (section 5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [150, 151]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make two key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [19]) and 2) outperforms an optimized state-of-the-art system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [19] refresh-access parallelization) by 4.1%. SMD provides 88.3% of the speedup provided by an oracle (No-Refresh) that completely mitigates the overhead of maintenance operations, on average across all evaluated high memory intensity four core workloads. Second, SMD reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than prior techniques that enable refresh-access parallelization (e.g., [19, 149]).
@@ 4.1 Overview of SMD
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section 6) in the physical interface of a DRAM chip. This pin is used for sending the ACT_NACK signal from the DRAM chip to the MC.
@@ 4.2 Region Locking Mechanism
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffer in a subarray. We design and evaluate SMD assuming a default lock region size of subarrays. We evaluate SMD’s sensitivity to lock region size (for a DRAM bank with 256 subarrays) and find that increasing the number of lock regions in a bank beyond 16 (i.e., having fewer than 16 subarrays in a lock region) yields diminishing returns (section 9.2).
@@ 4.3 Controlling an SMD Chip
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It can keep re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum (configurable) delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation. An enforced minimum delay of ARI is sufficient to guarantee forward progress for memory requests (section 4.4). The system designer can set the enforced minimum delay to a multiple of ARI to enforce a larger time period between two maintenance operations targeting the same region.
Setting the ACT Retry Interval (ARI). Setting ARI to a very low or high value relative to the expected duration of the maintenance operations can have a negative impact on system performance and energy efficiency. We empirically find that is a favorable configuration for the three maintenance mechanisms that we evaluate in this work (section 5).
@@
We estimate the overall latency based on the latency of RD (read) in a conventional DRAM chip. RD has a latency breakdown that resembles the latency breakdown of a . RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [141], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Bitvector (LRB) is (section 6). Given the relatively low complexity of the LRB compared to the datapath that is involved during RD, the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations (see section 8), we find that small latencies (e.g., ) have a negligible effect on system performance mainly because the number of rejected ACTs constitutes a small portion of all ACTs.
@@ 4.4 Forward Progress for Memory Requests
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (other than the one accessed by ACT_NACK’d memory request) prevents retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is exactly what a carefully designed maintenance mechanism should avoid. Instead, a carefully designed maintenance mechanism alternates between different lock regions as it performs new maintenance operations (e.g., periodic refresh section 5.1).
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, an implementation of SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for at least ARI after the end of a maintenance operation in a lock region (section 4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions.777An SMD chip may reject a memory request’s activate command. This means that the memory scheduler, based on its scheduling algorithm, has deemed that request to be the most important request to schedule. Therefore, it should be in the best interest of the scheduling algorithm to serve this memory request as soon as possible. The ARI interval allows the memory controller to schedule one or multiple other memory requests by leveraging lock-region-level parallelism. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the DDR4 standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI. To ensure that the memory controller retries the rejected memory request in due time, we set the number of allowed column commands (RD/WR) to an open row before an older request (to another row) is serviced (i.e., the Cap parameter [160]). Based on ARI=, we use a Cap of 7 in our evaluations.
@@ 4.5 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable (i.e., slightly asynchronous) from the perspective of the MC. However, the practically-observed performance overheads of this drawback are very low, as we demonstrate in section 8.
@@ 5 SMD Maintenance Mechanisms
We propose SMD-based maintenance mechanisms for three use cases. SMD is not limited to these three use cases and it can be used to support more operations in DRAM (as described in section 5.4 and section 5.5).
@@ 5.1 Use Case 1: DRAM Refresh
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of every tREFI interval . SMD-FR allows up to N refresh operations to be accumulated in the pending refresh counter. We set N=8.999DDR4 [141, 153, 158] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first [19, 36]. DDR5 reduces this number to [119]. Because the MC can keep a row open for a limited time (section 4.3), the pending refresh counter never exceeds the value N.

### v9

@@ 1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4, 5, 6]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrinks, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [7, 8, 9, 10, 11, 12, 13, 14, 2, 15, 5, 16, 17, 18]. A modern DRAM chip requires three major types of maintenance operations (described in detail in section˜5) for reliable and secure operation: 1) DRAM refresh [19, 20, 17, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], 2) RowHammer protection [47, 48, 49, 50, 12, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 18, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122], and 3) memory scrubbing [123, 124, 125, 126, 127, 128, 129, 130, 131, 17, 132, 133, 134, 135].111Memory scrubbing is not always required in consumer systems but often used in cloud systems. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [119, 136, 137]) and introducing new types of maintenance operations (e.g., targeted refresh [66, 64, 138], DDR5 RFM [119], and PRAC [119] as RowHammer defenses).222 A very recent update to the DDR5 standard [119] introduces PRAC, which is an on-DRAM-die read disturbance mitigation mechanism. PRAC requires more changes to the DRAM interface and continues to use RFM. Note that PRAC is concurrent with this work, as the initial version of this paper [139] was placed on arXiv on 27 July 2022 and initial submission to the MICRO 2022 conference was made on 22 April 2022.
@@
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.444SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors [4, 6] because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers). To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section˜4.5). While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations as SMD overlaps the latency of a maintenance operation in a DRAM region with memory accesses to other DRAM regions.
@@
To enable practical adoption of SMD, we propose to implement it with low-cost modifications to existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section˜6). Second, SMD implements a small (e.g., 16 bits per DRAM bank) memory structure called the Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [148, 149, 19]). In section˜6, we show that these modifications have low DRAM chip area (1.1% of a DRAM chip) and negligible latency overhead (only 0.4% increase in row activation latency).
@@
We demonstrate the usefulness and versatility of SMD by implementing three in-DRAM maintenance mechanisms for DRAM refresh (section˜5.1), RowHammer protection (section˜5.2), and memory scrubbing (section˜5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [150, 151]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make two key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [19]) and 2) outperforms an optimized state-of-the-art system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [19] refresh-access parallelization) by 4.1%. SMD provides 88.3% of the speedup provided by an oracle (No-Refresh) that completely mitigates the overhead of maintenance operations, on average across all evaluated high memory intensity four core workloads. Second, SMD reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than prior techniques that enable refresh-access parallelization (e.g., [19, 149]).
@@ 4.1 Overview of SMD
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section˜6) in the physical interface of a DRAM chip. This pin is used for sending the ACT_NACK signal from the DRAM chip to the MC.
@@ 4.2 Region Locking Mechanism
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffer in a subarray. We design and evaluate SMD assuming a default lock region size of subarrays. We evaluate SMD’s sensitivity to lock region size (for a DRAM bank with 256 subarrays) and find that increasing the number of lock regions in a bank beyond 16 (i.e., having fewer than 16 subarrays in a lock region) yields diminishing returns (section˜9.2).
@@ 4.3 Controlling an SMD Chip
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It can keep re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum (configurable) delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation. An enforced minimum delay of ARI is sufficient to guarantee forward progress for memory requests (section˜4.5). The system designer can set the enforced minimum delay to a multiple of ARI to enforce a larger time period between two maintenance operations targeting the same region.
Setting the ACT Retry Interval (ARI). Setting ARI to a very low or high value relative to the expected duration of the maintenance operations can have a negative impact on system performance and energy efficiency. We empirically find that is a favorable configuration for the three maintenance mechanisms that we evaluate in this work (section˜5).
@@
We estimate the overall latency based on the latency of RD (read) in a conventional DRAM chip. RD has a latency breakdown that resembles the latency breakdown of a . RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [141], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Bitvector (LRB) is (section˜6). Given the relatively low complexity of the LRB compared to the datapath that is involved during RD, the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations (see section˜8), we find that small latencies (e.g., ) have a negligible effect on system performance mainly because the number of rejected ACTs constitutes a small portion of all ACTs.
4.4 ACT_NACK Divergence Across DRAM Chips
SMD maintenance operations take place independently in each DRAM chip. Therefore, when a DRAM rank, which can be composed of multiple DRAM chips operating in lock step, receives an ACT command, some of the chips may send an ACT_NACK while others do not. Normally, ACT_NACK divergence would not happen for maintenance mechanisms that perform the exact same operation at the exact same time in all DRAM chips (e.g., the Fixed Refresh mechanism, Section 5.1). However, a mechanism can also operate differently in each DRAM chip, e.g., a variable rate refresh mechanism [20, 17, 41] or a RowHammer protection mechanism (e.g., [12, 73, 61, 119]). As a result of this divergence, the row becomes partially open: open in chips that do not send ACT_NACK and closed in chips that do.
The ACT_NACK divergence problem does not happen when each chip is individually controlled by the MC. As a solution for systems where a rank of chips operate in lock step, the MC can employ one of the three strategies: Precharge, Wait, and Hybrid. We evaluate each technique under common-case and worst-case scenarios with regard to when maintenance operations happen across different SMD chips in the same rank in section˜9.9.
Precharge. With the Precharge policy, the MC issues a PRE command to close the partially activated row when some DRAM chips send ACT_NACK but others do not. After closing the partially activated row, the MC can attempt to activate a row from a different lock region in the same bank.
Wait. With the Wait policy, the MC issues multiple ACT commands until a partially activated row becomes fully activated. When some chips send ACT_NACK for a particular ACT but others do not, the MC waits for ARI and issues a new ACT to attempt activating the same row in DRAM chips that previously sent ACT_NACK.
Hybrid. We also design a Hybrid policy, where the memory controller uses the Precharge policy to close a partially activated row if the request queue contains or more requests that need to access rows in different lock regions in the same bank. If the requests queue has less than requests to different lock regions, the MC uses the Wait policy to retry activating the rest of the partially activated row.
@@ 4.5 Forward Progress for Memory Requests
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (other than the one accessed by ACT_NACK’d memory request) prevents retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is exactly what a carefully designed maintenance mechanism should avoid. Instead, a carefully designed maintenance mechanism alternates between different lock regions as it performs new maintenance operations (e.g., periodic refresh section˜5.1).
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, an implementation of SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for at least ARI after the end of a maintenance operation in a lock region (section˜4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions.777An SMD chip may reject a memory request’s activate command. This means that the memory scheduler, based on its scheduling algorithm, has deemed that request to be the most important request to schedule. Therefore, it should be in the best interest of the scheduling algorithm to serve this memory request as soon as possible. The ARI interval allows the memory controller to schedule one or multiple other memory requests by leveraging lock-region-level parallelism. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the DDR4 standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI. To ensure that the memory controller retries the rejected memory request in due time, we set the number of allowed column commands (RD/WR) to an open row before an older request (to another row) is serviced (i.e., the Cap parameter [160]). Based on ARI=, we use a Cap of 7 in our evaluations.
@@ 4.6 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable (i.e., slightly asynchronous) from the perspective of the MC. However, the practically-observed performance overheads of this drawback are very low, as we demonstrate in section˜8.
@@ 5 SMD Maintenance Mechanisms
We propose SMD-based maintenance mechanisms for three use cases. SMD is not limited to these three use cases and it can be used to support more operations in DRAM (as described in section˜5.4 and section˜5.5).
@@ 5.1 Use Case 1: DRAM Refresh
SMD-FR (Fig. 4) operates independently in each DRAM bank and it uses three counters for managing the refresh operations: pending refresh counter, lock region counter, and row address counter. The pending refresh counter is initially zero and SMD-FR increments it by one at the end of every tREFI interval . SMD-FR allows up to N refresh operations to be accumulated in the pending refresh counter. We set N=8.999DDR4 [141, 153, 158] allows the MC to postpone issuing up to REF commands in order to serve pending memory requests first [19, 36]. DDR5 reduces this number to [119]. Because the MC can keep a row open for a limited time (section˜4.3), the pending refresh counter never exceeds the value N.